OpenStudy (anonymous):

Which measure is equivalent to 48 ft?

OpenStudy (_michaeld22):

Do you have any answer choices?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

hang on ill post

OpenStudy (anonymous):

A. 4 IN. B. 16 IN. C. 288 IN. D. 567 IN.

OpenStudy (_michaeld22):

1 ft. = 12 inches. 4/12 = .33ft 16/12 = 1.33ft 288/12 = 24ft 567/12 = 47.25ft The closest answer would be the final one.

OpenStudy (anonymous):

??....

OpenStudy (dannyrod2000):

12*48=567 in

OpenStudy (anonymous):

thanks

OpenStudy (dannyrod2000):

there are 12 inches in each feet so you multiply it by 48

OpenStudy (dannyrod2000):

to give you in inches

OpenStudy (_michaeld22):

Oh lol I did it backwards, didn't I? @dannyrod2000

jimthompson5910 (jim_thompson5910):

it looks like you wrote the digits of 576 in the wrong order

OpenStudy (dannyrod2000):

lol yup

jimthompson5910 (jim_thompson5910):

since it should be 48*12 = 576

OpenStudy (_michaeld22):

I was confused about that @jim_thompson5910 Thought it might've been a typo.

OpenStudy (anonymous):

yeah it was backwords thats why I was confussed